OpenStudy (anonymous):

Write the general conic form equation for a parabola with vertex at (2, 2) and focus at (1/2, 2). i got this y² - 4y + 2x = 0 Other choices y² - 4y + 6x - 8 = 0 y² - 4y - 6x + 16 = 0

OpenStudy (mertsj):

What is the focus?

OpenStudy (anonymous):

1/2,2

OpenStudy (mertsj):

x=-(y-k)^2+h x=-(y-2)^2+2 x=-y^2+4y-4+2 -y^2+4y-2-x=0 or y^2-4y+2+x=0

OpenStudy (anonymous):

thanks mertsj

OpenStudy (mertsj):

yw

OpenStudy (phi):

Let me suggest an alternate answer